\begin{thebibliography}{11}
\bibitem{aigner79}
M.~Aigner, 
{\sl Combinatorial Theory},
Grundlehren der Math. Wiss. {\bf 234}, Springer-Verlag, Berlin-New York, 1979.

\bibitem{barg97}
A.~Barg,
The matroid of supports of a linear code, 
{\sl Appl. Algebra Engrg. Commun. Comput.} {\bf 8} (1997), pp. 165--172.

\bibitem{britz10}
T.~Britz, 
Code enumerators and Tutte polynomials, 
{\sl IEEE Trans. Inform. Theory} {\bf 56} (2010), 4350--4358.

\bibitem{BS16}
T.~Britz and K.~Shiromoto,
On the covering dimension of a linear code,
{\sl IEEE Trans. Inform. Theory} {\bf 62} (2016), 2694--2701.

\bibitem{BSW15}
T.~Britz, K.~Shiromoto, and T.~Westerb\"ack, 
Demi-matroids from codes over finite Frobenius rings, 
{\sl Des. Codes Cryptogr.} {\bf 75} (2015), pp. 97--107.

\bibitem{CGB13}
T.H.~Chan, A.~Grant, and T.~Britz, 
Quasi-uniform codes and their applications, 
{\sl IEEE Trans. Inform. Theory} {\bf 59} (2013), 7915--7926.

\bibitem{CGLR18}
J.~Cruz, E.~Gorla, H.H.~L\'opez and A.~Ravagnani,
Weight distribution of rank-metric codes,
{\sl Des. Codes Cryptogr.} {\bf 86} (2018), pp. 1--16.

\bibitem{delsarte78}
P.~Delsarte, 
Bilinear forms over a finire field, with applications to coding theory,
{\sl Journal of Combinatorial Theory, Series A} {\bf 25} (1978), pp. 226--241.

\bibitem{gabidulin85a}
E.M.~Gabidulin, 
Theory of codes with maximum rank distance (Russian), 
{\sl Problemy Peredachi Informatsii}~{\bf 21}, 3--16 (1985).

\bibitem{gabidulin85b}
E.M.~Gabidulin,
Theory of codes with maximum rank distance (Translation),
{\sl Probl. Inf. Transm.}~{\bf 1}, 1--12 (1985).

\bibitem{GaPaTr91}
E.M.~Gabidulin, A.V.~Paramonov and O.V.~Tretjakov, 
Ideals over a non-commutative ring and their application in cryptology, 
in Proceedings of the Workshop on the Theory and Application of Cryptographic Techniques (EUROCRYPT ’91), vol. 547 of Lecture Notes in Computer Science, pp. 482--489, Brighton, UK, April 1991.

\bibitem{gadouleau2007}
M.~Gadouleau and Z.~Yan,
MacWilliams identity for the rank metric,
{\sl ISIT2007}, Nice, France, June 24 - June 29, 2007, pp. 36--40.

\bibitem{gadouleau2008}
M.~Gadouleau and Z.~Yan,
MacWilliams identity for the rank metric,
{\sl EURASIP Journal on Wireless Communications and Networking},
vol. {\bf 2008}, Article ID 754021, 13 pages.

\bibitem{greene76}
C.~Greene,
Weight enumeration and the geometry of linear codes, 
{\sl Stud. Appl. Math.} {\bf 55} (1976), pp. 119--128.

\bibitem{huffman2003}
W.~Huffman and V.~Pless,
{\sl Fundamentals of Error-Correcting Codes},
Cambridge University Press, Cambridge, 2003.

\bibitem{KMS17}
Y.~Koga, T.~Maruta and K.~Shiromoto,
On critical exponents of Dowling matroids,
{\sl Des. Codes Cryptogr.} (2017). https://doi.org/10.1007/s10623-017-0431-8

\bibitem{KsGa05}
A.~Kshevetskiy and E.N.~Gabidulin,
The new construction of rank codes,
{\sl Proceedings of IEEE International Symposium on Information Theory} (ISIT05), pp. 2105--2108, Adelaide, Australia, 2005.

\bibitem{lam99}
T.Y.~Lam, 
{\sl Lectures on modules and rings}, 
Graduate Texts in Mathematics, vol. {\bf 189}, Springer-Verlag, New York, 1999.
 
\bibitem{MS77}
F.~J.~MacWilliams and N.~J.~A.~Sloane,
{\sl The theory of error-correcting codes}, North-Holland, Amsterdam 1977.

\bibitem{oxley92}
J.~Oxley,
{\sl Matroid Theory},
Oxford University Press, New York, 1992.

\bibitem{Oxley-Whittle93}
J.~Oxley and G.~Whittle,
A characterization of Tutte invariants of $2$-polymatroids,
{\sl Journal of Combinatorial Theory, Series B} {\bf 59} (1993), pp. 210--244.

\bibitem{ravagnani2016}
A.~Ravagnani, 
Rank-metric codes and their duality theory,
{\sl Designs, Codes, and Cryptography} {\bf 80} (2016), pp. 197--216.

\bibitem{shiromoto99}
K.~Shiromoto,
A basic exact sequence for the Lee and Euclidean weights of linear codes over ${\mathbb Z}_{\ell}$,
{\sl Linear Algebra and its Applications} {\bf 295} (1999), pp. 191--200.

\bibitem{SiKsKo08}
D.~Silva, F.R.~Kschischang and R.~Koetter,
A rank-metric approach to error control in random network coding, 
{\sl IEEE Trans. Inform. Theory}~{\bf 54}, 3951--3967 (2008).

\bibitem{simonis94}
J.~Simonis, 
The effective length of subcodes,
{\sl Appl. Algebra Engrg. Commun. Comput.} {\bf 5} (1994), pp. 371--377.

\bibitem{simonis98}
J.~Simonis and A.~Ashikhmin, 
Almost affine codes, 
{\sl Des. Codes Cryptogr.}, {\bf 14} (1998), pp. 179--197.

\bibitem{TaSeCa98}
V.~Tarokh, N.~Seshadri and A.R.~Calderbank,
Space-time codes for high data rate wireless communication: performance criterion and code construction, 
{\sl IEEE Trans. Inform. Theory}~{\bf 44}, 744--765 (1998).

\bibitem{vertigan04}
D.~Vertigan, 
Latroids and their representation by codes over modules,
{\sl Trans. Amer. Math. Soc.} {\bf 356} (2004), pp. 3841--3868.

\bibitem{welsh76}
D.J.A.~Welsh,
{\sl Matroid Theory},
Academic Press, London, 1976.
\end{thebibliography}

\documentclass[a4paper, 12pt]{article}
\usepackage{latexsym,amsmath,amssymb,bm, color}
\usepackage{amsmath,amssymb,color}

\newtheorem{thm}{Theorem}
\newtheorem{prop}[thm]{Proposition}
\newtheorem{lem}[thm]{Lemma}
\newtheorem{defi}[thm]{Definition}

\newcommand{\qed}{\hbox{\rule[-2pt]{3pt}{6pt}}}

\newcommand{\pf}{{\bf Proof. \ }}
\renewcommand{\qed}{\hfill $\Box$

\medskip}

\begin{document}
\title{\bf Matroids and Codes with the Rank Metric} 
\author{Keisuke Shiromoto\footnote{This work was supported by JSPS KAKENHI Grant Number 17K05348.}\\
Department of Mathematics and Engineering, \\
Kumamoto University, \\
2-39-1, Kurokami, Kumamoto 860-8555, Japan\\
           \tt{keisuke@kumamoto-u.ac.jp}
           } 
\maketitle
\begin{abstract}
We study the relationship between a $q$-analogue of matroids and linear codes with the rank metric in the vector space of matrices with entries in a finite field. We prove a Greene type identity for the rank generating function of these matroidal structures and the rank weight enumerator of these linear codes.
As an application, we give a combinatorial proof of a MacWilliams type identity for Delsarte rank-metric codes.\\

\noindent
{\em Keywords:} rank-metric code; matroid; polymatroid; Greene's identity; MacWilliams identity\\

\noindent
{\em Mathematics Subject Classification:}
05B35, 94B60

\end{abstract}

\section{Introduction}
Let $S$ be a finite set and let $\rho \: : \: 2^{S}\rightarrow {\mathbb Z}$ be a function.
An ordered pair $M=(S, \rho)$ is called a {\em matroid} if
\begin{enumerate}
\item[{\rm (1)}] If $X \subseteq S$, then $0 \leq \rho(X) \leq |X|$.
\item[{\rm (2)}] If $X,Y\subseteq S$ and $X \subseteq Y$, then $\rho(X) \leq \rho(Y)$.
\item[{\rm (3)}] If $X,Y\subseteq S$, then $\rho(X \cup Y)+\rho(X \cap Y) \leq \rho(X)+\rho(Y).
$
\end{enumerate}
Matroids are constructed from a lot of combinatorial or algebraic structures including graphs and matrices (cf. \cite{welsh76, oxley92}).
Let ${\mathbb F}_q$ be a finite filed of $q$ elements.
For an $[n,k]$ code $C$ over ${\mathbb F}_q$ and any subset $X \subseteq S:=\{1,2,\ldots,n\}$, define
$$
\rho(X):=\dim C\setminus (S-X) (=\dim C-\dim C(S-X)),
$$
where, for any  $Y \subseteq S$, $C\setminus Y$ denotes the punctured code by $Y$ and $$C(Y):=\{{\bm x} \in C \: : \: \mbox{\rm supp}({\bm x})\subseteq Y\}.$$
Then we find that $M_C:=(S,\rho)$ satisfies the above conditions and so $M_C$ is a matroid.

The (Whitney) {\em rank generating function} of a matroid $M=(S, \rho)$ is defined by
$$
R(M;x,y):=\sum_{X \subseteq S}x^{\rho(S)-\rho(X)}y^{|X|-\rho(X)}.
$$
In 1976, Greene (\cite{greene76}) proved one celebrated result, known as Greene's identity, on the relationship between the Hamming weight enumerator $W_C(x,y)$ of an $[n,k]$ code $C$ over ${\mathbb F}_q$ and the rank generating function of the corresponding matroid $M_C$ as follows:
$$
W_C(x,y)=y^{n-\dim C}(x-y)^{\dim C}R(M_C;\frac{qy}{x-y}, \frac{x-y}{y}).
$$
As an application of this result, he gave a elegant combinatorial proof of the MacWilliams identity for the Hamming weight enumerator (cf. \cite{welsh76}).
After this paper, a number of paper in the link on matroids and codes have been published, for instance, generalizations of the identity to higher weight enumerators (\cite{barg97, britz10, simonis94}), and to some enumerators for other classes of codes (\cite{BSW15, CGB13, simonis98, vertigan04}). 
In particular, Britz (\cite{britz10}) proved the equivalence between the set of higher weight enumerators of a linear code over a finite filed and the rank generating function associated to the code.
In \cite{BS16} and \cite{KMS17}, the authors studied the critical problem, known as a classical problem in matroid theory, from coding theoretical approach.

Delsarte introduced a rank-metric code as a set of matrices of given size over a finite field in \cite{delsarte78}.
He mainly studied the rank-metric code based on his research on association schemes.
In \cite{gabidulin85a, gabidulin85b}, Gabidulin gave a different definition of rank-metric codes such as a set of vectors in a vector space over an extension field.
The rank-weight of a codeword is defined by the rank of the associated matrix.
Ravagnani proved that Gabidulin codes can be regarded as a special case of Delsarte codes and he compared the duality theories of these codes in \cite{ravagnani2016}.
Some useful applications of rank-metric codes are widely known, for instance, to space-time codes (\cite{TaSeCa98}), to network coding (\cite{SiKsKo08}), and to cryptography (\cite{GaPaTr91}).

In addition, some of the results in classical coding theory are generalized to rank-metric codes. 
For instance, a Singleton bound on the minimum rank distance of the codes are proven in \cite{delsarte78}.
Some constructions of the codes which attain the bound, called {\em maximum rank distance} (MRD) {\em codes}, are known (\cite{CGLR18, delsarte78, gabidulin85a, gabidulin85b, KsGa05}).
The MacWilliams type identitis for rank weight enumerators of rank-metric codes are proposed as several expressions: an algebraic expression (\cite{delsarte78}), identitis for rank weight enumerators of Gabidulin rank-metric codes closed to the original form (\cite{gadouleau2007, gadouleau2008}), and moments of rank distributions of Delsarte rank-metric codes (\cite{ravagnani2016}).

The layout of this paper is as follows.
In Section 2, we firstly introduce some basic notion on Delsarte rank-metric codes and new matroidal structures related rank-metric codes.
And we give a Greene type identity for rank-metric codes.
Section 3 contains an application of the Greene type identity to the MacWilliams identity for Delsarte rank-metric codes.

Most of the basic terminology in coding theory and matroid theory used in this paper will be found in standard texts (cf. \cite{huffman2003, MS77, oxley92, welsh76}).

\section{$(q, r)$-Polymatroids and Rank Generating Functions}

We denote by $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$ the ${\mathbb F}_q$-vector space of $n \times m$ matrices with entries in ${\mathbb F}_q$.
A {\em Delsarte rank-metric code} ${\cal C}$ of size $n \times m$ over ${\mathbb F}_q$ is an ${\mathbb F}_q$-linear subspace of $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
Throughout this paper, set $E:={\mathbb F}_q^n$.
For convenience, we shall write for $D \leq E$ if $D$ is a subspace of $E$.
For any subspace $J \leq E$, define
\begin{eqnarray*}
{\cal C}(J) &:=&\{M \in {\cal C}\: |\: \mbox{\rm col}(M) \subseteq J\},\\
J^{\perp}&:=&\{{\bm y} \in {\mathbb F}_q^n \: | \: {\bm x} \cdot {\bm y}=0,\: {}^{\forall}\!{\bm x } \in J\},
\end{eqnarray*}
where $\mbox{\rm col}(M)$ denotes the column space of $M$ over ${\mathbb F}_q$.

\begin{lem}
\label{lem:subspaces}
${\cal C}(J)$ is an ${\mathbb F}_q$-linear subspace of $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
\end{lem}
\pf
For any $M, N \in {\cal C}(J)$, let ${\bm u}_1,\ldots, {\bm u}_m$ and ${\bm v}_1,\ldots, {\bm v}_m$ be the column vectors of $M$ and $N$, respectively.
For any $\alpha, \beta \in {\mathbb F}_q$, take a vector ${\bm x} \in \mbox{\rm col}(\alpha M+\beta N)$.
Then ${\bm x}$ can be written as
\begin{eqnarray*}
{\bm x}&=&a_1(\alpha {\bm u}_1+\beta {\bm v}_1)+\cdots+a_m(\alpha {\bm u}_m+\beta {\bm v}_m)\\
&=&\alpha(a_1 {\bm u}_1+\cdots+a_m {\bm u}_m)+\beta(a_1 {\bm v}_1+\cdots+a_m {\bm v}_m),
\end{eqnarray*}
for some $a_1,\ldots, a_m \in {\mathbb F}_q$.
Since $\alpha(a_1 {\bm u}_1+\cdots+a_m {\bm u}_m) \in J$ and $\beta(a_1 {\bm v}_1+\cdots+a_m {\bm v}_m) \in J$, we have that ${\bm x} \in J$.
\qed

We denote by $\Sigma(E)$ the set of all subspaces of $E$ and denote by $\Sigma(J)$ the set of all subspaces of any subspace $J \in \Sigma(E)$.

Now we shall introduce a $(q,r)$-polymatroid as a $q$-analogue of $k$-polymatroids (cf. \cite{Oxley-Whittle93}).
\begin{defi}
{\rm 
A $(q, r)$-{\em polymatroid} is an ordered pair $P=(E, \rho)$ consisting of a vector space $E:={\mathbb F}_q^n$ and a function $\rho \: : \: \Sigma(E)\to {\mathbb Z}^{+}\cup \{0\}$ having the following properties:
\begin{enumerate}
\item[{\rm (R1)}] If $A \leq E$, then $0 \leq \rho(A) \leq r \dim A$.
\item[{\rm (R2)}] If $A, B \leq E$ and $A \subseteq B$, then $\rho(A) \leq \rho(B)$.
\item[{\rm (R3)}] If $A, B \leq E$, then $\rho(A+B)+\rho(A \cap B) \leq \rho(A)+\rho(B).
$
\end{enumerate}
}
\end{defi}

We find easily that $\rho(\{{\bm 0}\})=0$ from (R1) and a $(q,1)$-polymatroid is a $q$-analogue of matroids.

\begin{prop}
\label{code-q-matroid}
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$, and let $\rho\: : \: \Sigma(E) \to {\mathbb Z}^{+}\cup \{0\}$ be the function defined by 
$$
\rho(J):=\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}(J^{\perp}),
$$
for any subspace $J$ of $E$.
Then $P_{\cal C}:=(E, \rho)$ is a $(q, m)$-polymatroid.
\end{prop}
\pf
Clearly the function $\rho$ satisfies (R1) and (R2).
Now we prove that (R3) holds.
Let $J_1$ and $J_2$ are subspaces of $E$.
From Lemma \ref{lem:subspaces}, we have that
$$
\dim {\cal C}(J_1^{\perp})+\dim {\cal C}(J_2^{\perp})=\dim {\cal C}(J_1^{\perp})\cap {\cal C}(J_2^{\perp})+\dim({\cal C}(J_1^{\perp})+{\cal C}(J_2^{\perp})).
$$
Moreover, it follows that
\begin{eqnarray*}
&& {\cal C}(J_1^{\perp})\cap {\cal C}(J_2^{\perp})={\cal C}(J_1^{\perp} \cap J_2^{\perp})={\cal C}((J_1+J_2)^{\perp}),\\
&& {\cal C}(J_1^{\perp})+ {\cal C}(J_2^{\perp}) \subseteq {\cal C}(J_1^{\perp}+J_2^{\perp})={\cal C}((J_1 \cap J_2)^{\perp}),
\end{eqnarray*}
because of $\mbox{\rm col}(M_1+M_2) \subseteq \mbox{\rm col}(M_1)+\mbox{\rm col}(M_2)$ for any $M_1, M_2 \in \mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
Therefore we have that
\begin{eqnarray*}
\rho(J_1)+\rho(J_2)&=&(\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}(J_1^{\perp}))+(\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}(J_2^{\perp}))\\
&\geq&(\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}((J_1+J_2)^{\perp}))
+(\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}((J_1\cap J_2)^{\perp}))\\
&=&\rho(J_1+J_2)+\rho(J_1\cap J_2).
\end{eqnarray*}
Thus $P_{{\cal C}}=(E, \rho)$ is a $(q, m)$-polymatroid.
\qed

The {\em rank generating function} of  a $(q, r)$-polymatroid $P=(E, \rho)$ is defined by
$$
R_{P}(X_1, X_2, X_3, X_4):=\sum_{D \in \Sigma(E)}f_{P}^D(X_1, X_2)g^{\dim D}(X_3, X_4),
$$
where
$$
f_{P}^J(X, Y):=X^{\rho(E)-\rho(J)}Y^{r\dim J-\rho(J)},
$$
for any subspace $J \in \Sigma(E)$, and
$$
g^{l}(X, Y):=\prod_{i=0}^{l-1}(X-q^iY),
$$
for any nonnegative integer $l \in {\mathbb Z}^{+}\cup\{0\}$.

The following lemma is essential.
\begin{lem}
\label{lem:rank}
Let $P=(E,\rho)$ be a $(q, r)$-polymatroid.
If $A$ and $B$ are subspaces of $E$ such as $A \subseteq B$, then
$$
0 \leq \rho(A)-\rho(B)-r(\dim A-\dim B).
$$
\end{lem}
\pf
Let $D$ be a subspace of $E$ generated by $B-A$.
From (R1), we have that
$$
\rho(D) \leq r\dim D=r(\dim B-\dim A).
$$
From (R3), we also have that
$$
\rho(A)+\rho(D)\geq \rho(A+D)+\rho(A\cap D)\geq\rho(B).
$$
By combining the above two inequalities, the lemma follows.
\qed

\begin{prop}
\label{dual}
For any $(q, r)$-polymatroid $P=(E, \rho)$ and any subspace $J$ of $E$, define
$$
\rho^*(J):=\rho(J^{\perp})+r \dim J-\rho(E).
$$
Then $P^*=(E, \rho^*)$ is a $(q, r)$-polymatroid.
\end{prop}
\pf
$J^{\perp} \subseteq E$ implies that $\rho(J^{\perp}) \leq \rho(E)$.
So we have that $(0\leq)\rho^*(J)\leq r \dim J$.
Thus (R1) holds.

Let $J_1$ and $J_2$ are subspaces of $E$.
Assume that $J_1 \subseteq J_2$. 
Then it follows from Lemma \ref{lem:rank} that
\begin{eqnarray*}
\rho^*(J_2)-\rho^*(J_1)&=&(\rho(J_2^{\perp})+r \dim J_2-\rho(E))-(\rho(J_1^{\perp})+r \dim J_1-\rho(E))\\
&=&\rho(J_2^{\perp})-\rho(J_1^{\perp})+r(\dim J_2-\dim J_1)\\
&=&\rho(J_2^{\perp})-\rho(J_1^{\perp})+r(\dim J_1^{\perp}-\dim J_2^{\perp})\\
&\geq&0.
\end{eqnarray*}
Hence (R2) holds.

From the definition of $\rho^*$, it follows that
\begin{eqnarray*}
\rho^*(J_1\cap J_2)&=&\rho((J_1\cap J_2)^{\perp})+r\dim (J_1\cap J_2)-\rho(E),\\
\rho^*(J_1+ J_2)&=&\rho((J_1+ J_2)^{\perp})+r\dim (J_1+ J_2)-\rho(E),\\
\rho^*(J_1)&=&\rho(J_1^{\perp})+r\dim J_1-\rho(E),\\
\rho^*(J_2)&=&\rho(J_2^{\perp})+r\dim J_2-\rho(E).
\end{eqnarray*}
From the inequality in (R3) for $\rho$, we have that
\begin{eqnarray*}
&&\rho^*(J_1\cap J_2)+\rho^*(J_1+ J_2)-\rho^*(J_1)-\rho^*(J_2)\\
&=&\{\rho(J_1^{\perp}+ J_2^{\perp})+\rho(J_1^{\perp}\cap J_2^{\perp})
 -(\rho(J_1^{\perp})+\rho(J_2^{\perp}))\}\\
 &&\hfill+r\{\dim (J_1\cap J_2)+\dim (J_1+ J_2)-(\dim J_1+\dim J_2)\}\\
&\leq& 0.
\end{eqnarray*}
Then (R3) holds.
\qed

For a rank generating function $R_{P}(X_1, X_2, X_3, X_4)$ of  a $(q, r)$-polymatroid $P=(E, \rho)$, define
$$
\widehat{R}_{P}(X_1, X_2, X_3, X_4):=\sum_{D \in \Sigma(E)}f_{P}^D(X_1, X_2)g^{\dim D^{\perp}}(X_3, X_4).
$$
The following result is a duality on rank generating functions of $(q, r)$-polymatroids.

\begin{thm}
\label{thm:duality}
Let $P=(E, \rho)$ be a $(q, r)$-polymatroid.
Then
$$
R_{P^*}(X_1, X_2, X_3, X_4)=\widehat{R}_{P}(X_2, X_1, X_3, X_4).
$$
\end{thm}
\pf
It follows from the definition of rank generating functions that
\begin{eqnarray*}
R_{P^*}(X_1, X_2, X_3, X_4)&=&\sum_{D \in \Sigma(E)}f_{P^*}^D(X_1, X_2)g^{\dim D}(X_3, X_4)\\
&=&\sum_{D \in \Sigma(E)}X_1^{\rho^*(E)-\rho^*(D)}X_2^{r\dim D-\rho^*(D)}
\prod_{i=0}^{\dim D-1}(X_3-q^iX_4)\\
&=&\sum_{D \in \Sigma(E)}X_1^{(rn-\rho(E))-(\rho(D^{\perp})+r\dim D-\rho(E))}X_2^{r\dim D-(\rho(D^{\perp})+r\dim D-\rho(E))}\\
&& \hspace{8cm} \times \prod_{i=0}^{\dim D-1}(X_3-q^iX_4)\\
&=&\sum_{D \in \Sigma(E)}X_1^{r(n-\dim D)-\rho(D^{\perp})}X_2^{\rho(E)-\rho(D^{\perp})}
\prod_{i=0}^{\dim D-1}(X_3-q^iX_4)\\
&=&\sum_{D' \in \Sigma(E)}X_2^{\rho(E)-\rho(D')}X_1^{r\dim D'-\rho(D')}
\prod_{i=0}^{n-\dim D'-1}(X_3-q^iX_4)\\
&=&\widehat{R}_{P}(X_2, X_1, X_3, X_4).
\end{eqnarray*}
\qed

The {\em trace product} of matrices $M, N \in \mbox{\rm Mat}(n\times m, {\mathbb F}_q)$
is defined by
$$
\langle M, N\rangle :=\mbox{\rm Tr}(MN^{{\mbox{\rm T}}}),
$$
where $\mbox{\rm Tr}$ denotes the trace of a matrix.
Then it is easy to find that the trace product can be calculate from the standard inner products of each row vectors as follows:

\begin{lem}
\label{lem:trace-products}
Let $M_i$ and $N_i$ be the $i$-th column vectors of $M$ and $N$, respectively.
Then we have that
$$
\langle M, N\rangle=\sum_{i=1}^m M_i\cdot N_i,
$$
where $M_i\cdot N_i$ denotes the standard inner product of vectors $M_i$ and $N_i$.
\end{lem}

Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
The {\em dual code} of ${\cal C}$ is defined by
$$
{\cal C}^{\perp}:=\{N \in \mbox{\rm Mat}(n\times m, {\mathbb F}_q)\: : \: \mbox{\rm $\langle M, N\rangle=0$ for all $M \in {\cal C}$}\}.
$$
Then the following result is well-known (see, Lemma 5 in \cite{ravagnani2016}).

\begin{lem} 
\label{lem:dual-codes}
{\rm (\cite{ravagnani2016})}
For any Delsarte rank-metric codes ${\cal C}, {\cal D} \in \mbox{\rm Mat}(n\times m, {\mathbb F}_q)$,
it follows that
\begin{enumerate}
\item[{\rm (1)}] $({\cal C}^{\perp})^{\perp}={\cal C}$;
\item[{\rm (2)}] $\dim _{{\mathbb F}_q} ({\cal C}^{\perp})=nm-\dim _{{\mathbb F}_q} ({\cal C})$;
\item[{\rm (3)}] $({\cal C} \cap {\cal D})^{\perp}={\cal C}^{\perp}+{\cal D}^{\perp}$, and $({\cal C} + {\cal D})^{\perp}={\cal C}^{\perp}\cap{\cal D}^{\perp}$.
\end{enumerate}
\end{lem}

Set ${\cal V}:=\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
Then it turns out easily that, for any subspace $R$ of ${\mathbb F}_q^n$,
\begin{enumerate}
\item[(1)] ${\cal C}(R)={\cal V}(R)\cap {\cal C}$, for any Delsarte rank-metric code ${\cal C} \subseteq {\cal V}$;
\item[(2)] $|{\cal V}(R)|=|R|^m=q^{m\dim R}$.
\end{enumerate}

\begin{lem}
\label{lem: duality}
For any subspace $R$ of ${\mathbb F}_q^n$, 
$$
({\cal V}(R))^{\perp}={\cal V}(R^{\perp}).
$$
\end{lem}
\pf
For any matrix $M\in {\cal V}(R^{\perp})$, Lemma \ref{lem:trace-products} implies that
\begin{eqnarray*}
\langle M, N\rangle &=&\mbox{\rm Tr}(MN^{{\mbox{\rm T}}})\\
&=&\sum_{i=1}^m M_i\cdot N_i=0,
\end{eqnarray*}
for all $N \in {\cal V}(R)$.
Thus we have that $M \in ({\cal V}(R))^{\perp}$ and so 
${\cal V}(R^{\perp}) \subseteq ({\cal V}(R))^{\perp}$.

Conversely, take a matrix $M \in  ({\cal V}(R))^{\perp}$.
Set
$$
R=\{{\bm x}_1,\ldots,{\bm x}_{|R|}\}.
$$
Let $M({\bm x}_i, j)$ be the $n \times m$ matrix over ${\mathbb F}_q$ with ${\bm x}_i$ as the $j$-th column vector and ${\bm 0}$ elsewhere.
Then it turns out that $M({\bm x}_i, j) \in {\cal V}(R)$.
Thus we have that
\begin{eqnarray*}
0&=& \langle M, M({\bm x}_i, j)\rangle=\sum_{l=1}^m M_l\cdot (M({\bm x}_i, j))_l\\
&=&M_j \cdot {\bm x}_i.
\end{eqnarray*}
This implies that $M_j \in R^{\perp}$, for all $j=1,2,\ldots, m$.
Therefore it follows that $M \in {\cal V}(R^{\perp})$ and so ${\cal V}(R^{\perp}) \supseteq ({\cal V}(R))^{\perp}$.
\qed

Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
The {\em dual space} ${\cal C}^{\star}$ of ${\cal C}$ is defined by
$$
{\cal C}^{\star}:=\mbox{\rm Hom}_{{\mathbb F}_q}({\cal C}, {\mathbb F}_q).
$$
Since ${\cal C}$ is a finite free ${\mathbb F}_q$-module, there is a (non-natural) isomorphism ${\cal C} \cong {\cal C}^{\star}$ (see, for instance, \cite{lam99}).

The following exact sequence is an analogue of the basic exact sequence on Proposition 2 in \cite{shiromoto99}.
 
\begin{prop}
\label{thm:basic-exact-sequence}
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
For any subspace $R$ of ${\mathbb F}_q^n$, there is an exact sequence as ${\mathbb F}_q$-modules,
$$
0 \longrightarrow {\cal C}^{\perp}(R) \xrightarrow{\scriptsize {\mbox{\rm inc}}} {\cal V}(R) \xrightarrow{ f }
{\cal C}^{\star} \xrightarrow{\scriptsize {\mbox{\rm res}}} {\cal C}(R^{\perp})^{\star} \longrightarrow 0,
$$
where $f$ is an ${\mathbb F}_q$-homomorphism defined by
\begin{eqnarray*}
f\: : \: {\cal V} &\longrightarrow& {\cal C}^{\star}\\
M &\longmapsto& (\hat{M}\: : \: N \longmapsto \langle M, N\rangle),
\end{eqnarray*}
and the maps {\rm inc} and {\rm res} denote the inclusion map and the restriction map, respectively.
\end{prop}
\pf
The inclusion map {\rm inc} is a natural injection, and the restriction map {\rm res} is surjective, since 
${\mathbb F}_q$ is an injective module over itself.

Take $M \in {\cal C}^{\perp}(R)$.
Then it follows that
$$
\hat{M}\: : \: N \longmapsto \langle M, N\rangle=0,
$$
for all $N \in {\cal C}$.
This implies that $M \in \ker f$ and so $\mbox{\rm Im}(\mbox{\rm inc})\subseteq \ker f$.
Conversely, take $M \in \ker f$.
Then we have that $\hat{M}(N)=\langle M, N\rangle=0$ for all $N \in {\cal C}$ and so $M \in {\cal C}^{\perp} \cap {\cal V}(R)={\cal C}^{\perp}(R)$.
Therefore we have that $\mbox{\rm Im}(\mbox{\rm inc})\supseteq \ker f$.
Thus it follows that $\mbox{\rm Im}(\mbox{\rm inc})= \ker f$.

Next we prove that $\ker (\mbox{\rm res})= \mbox{\rm Im} f $.
Take $M \in {\cal V}(R)$.
Then it also follows from Lemma \ref{lem:trace-products} that $\hat{M}(N)=\langle M, N\rangle=0$ for all $N \in {\cal C}(R^{\perp})$.
This implies that $\hat{M} \in \ker (\mbox{\rm res})$ and so $\mbox{\rm Im} f \subseteq \ker (\mbox{\rm res})$.
Conversely, take $\lambda \in \ker (\mbox{\rm res})$.
The map $f$ is surjective.
Thus there exists $M \in {\cal V}$ such that $\lambda=\hat{M}$.
Therefore we have that $\hat{M}(N)=\langle M, N\rangle=0$ for all $N \in {\cal C}(R^{\perp})$, and so $M \in ({\cal C}(R^{\perp}))^{\perp}$.
From Lemma \ref{lem: duality}, we have that
\begin{eqnarray*}
({\cal C}(R^{\perp}))^{\perp}&=&({\cal C}\cap {\cal V}(R^{\perp}))^{\perp}={\cal C}^{\perp}+{\cal V}(R^{\perp})^{\perp}\\
&=&{\cal C}^{\perp}+{\cal V}(R),
\end{eqnarray*}
and so $M \in {\cal C}^{\perp}+{\cal V}(R)$.
Then $M=M_1+M_2$ for some $M_1 \in {\cal C}^{\perp}$ and $M_2 \in {\cal V}(R)$.
Since $\hat{M_1}(N)=0$ for all $N \in {\cal C}$, it turns out that 
$\lambda=\hat{M}=\hat{M_2} \in \mbox{\rm Im} f$.
Thus it follows that $\mbox{\rm Im} f= \ker (\mbox{\rm res})$.

Therefore the proposition follows.
\qed

\begin{prop}
\label{prop:duality}
For any Delsarte rank-metric code ${\cal C}$ in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$, let $P_{\cal C}=(E, \rho)$ be the $(q, m)$-polymatroid obtained from ${\cal C}$ such as discussed in Proposition {\rm \ref{code-q-matroid}}.
Then $P_{\cal C}^{*}=P_{{\cal C}^{\perp}}$.
\end{prop}
\pf
Take any subspace $J \subseteq E$.
Set
$$
\tau(J):=\dim_{{\mathbb F}_q} {\cal C}^{\perp} -\dim_{{\mathbb F}_q} {\cal C}^{\perp}(J^{\perp}).
$$
From the definition of dual $(q, m)$-polymatroids, we have that
\begin{eqnarray*}
\rho^{*}(J)&=&\rho(J^{\perp})+m \dim J-\rho(E)\\
&=&(\dim_{{\mathbb F}_q} {\cal C}-\dim_{{\mathbb F}_q} {\cal C}(J))+m \dim J-\dim_{{\mathbb F}_q} {\cal C}\\
&=& m \dim J -\dim_{{\mathbb F}_q} {\cal C}(J).
\end{eqnarray*}
By applying $J^{\perp}$ for $R$ in Proposition \ref{thm:basic-exact-sequence}, we have that
\begin{eqnarray*}
&&|{\cal C}^{\perp}(J^{\perp})|\cdot |{\cal C}^{\star}|=|{\cal V}(J^{\perp})|\cdot |{\cal C}(J)^{\star}|\\
&\Longleftrightarrow& |{\cal C}^{\perp}(J^{\perp})|\cdot |{\cal C}|=|{\cal V}(J^{\perp})|\cdot |{\cal C}(J)|\\
&\Longleftrightarrow& q^{\dim_{{\mathbb F}_q} {\cal C}^{\perp}(J^{\perp})} \times q^{\dim_{{\mathbb F}_q} {\cal C}}=q^{m \dim J^{\perp}} \times q^{\dim_{{\mathbb F}_q} {\cal C}(J)}\\
&\Longleftrightarrow& \dim_{{\mathbb F}_q} {\cal C}^{\perp}(J^{\perp})+\dim_{{\mathbb F}_q} {\cal C}=m \dim J^{\perp}+\dim_{{\mathbb F}_q} {\cal C}(J)
\end{eqnarray*}
From the above equation and Lemma \ref{lem:dual-codes}(2), it follows that
\begin{eqnarray*}
\tau(J)&=& (nm-\dim_{{\mathbb F}_q} {\cal C})-(m \dim J^{\perp}+\dim_{{\mathbb F}_q} {\cal C}(J)-\dim_{{\mathbb F}_q} {\cal C})\\
&=& m \dim J-\dim_{{\mathbb F}_q} {\cal C}(J)\\
&=& \rho^{*}(J).
\end{eqnarray*}
\qed

\begin{defi}
{\rm (cf. \cite{gadouleau2008, ravagnani2016})
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
The {\em rank distribution} of ${\cal C}$ is the collection $\{A_i({\cal C})\}_{i\in {\mathbb Z}_{\geq 0}}$, where
$$
A_i({\cal C}):=|\{M \in {\cal C} \: : \: \mbox{\rm rank}(M)=i\}|.
$$
The {\em rank weight enumerator} of ${\cal C}$ is defined by
$$
W_{{\cal C}}^{{\rm R}} (x, y):=\sum_{i=0}^n A_i({\cal C}) x^{n-i}y^i.
$$
}
\end{defi}

For any Delsarte rank-metric code ${\cal C}$ in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$ and any subspace $R \subseteq  {\mathbb F}_q^n$, set
\begin{eqnarray*}
A_{{\cal C}}(R)&:=&|\{M \in {\cal C} \: : \: \mbox{\rm col}(M)=R\}|,\\
B_{{\cal C}}(R)&:=&|\{M \in {\cal C} \: : \: \mbox{\rm col}(M)\subseteq R\}|.
\end{eqnarray*}
Then it turns out easily that $B_{{\cal C}}(R)=\sum_{T \in \Sigma(R)}A_{{\cal C}}(T)$ and $B_{{\cal C}}(R)=|{\cal C}(R)|$.
From the M\"obius inversion formula (see, for instance, \cite{aigner79}), we have the following equation.

\begin{lem}
\label{lem:Mobius-inversion}
$$
A_{{\cal C}}(R)=\sum_{T \in \Sigma(R)}(-1)^{\dim R-\dim T} q^{{\dim R - \dim T \choose 2}}B_{{\cal C}}(T),
$$
where ${a \choose b}$ denotes the binomial coefficient of integers $a \geq b \geq 0$.
\end{lem}

Now we prove our main result in this paper.
The following equation is a kind of Greene type identities for $(q, r)$-polymatroids and Delsarte rank-metric codes.

\begin{thm}
\label{thm:greene-type-id}
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$, and let $P_{\cal C}=(E, \rho)$ be the $(q, m)$-polymatroid obtained from ${\cal C}$ such as discussed in Proposition {\rm \ref{code-q-matroid}}.
Then 
$$
W_{{\cal C}}^{{\rm R}} (x, y)=y^{n-\dim {\cal C} / m}
R_{P_{\cal C}}(qy^{1/m}, \frac{1}{y^{1/m}}, x, y).
$$
\end{thm}
\pf
By Lemma \ref{lem:Mobius-inversion} and the $q$-binomial Theorem, we obtain that
\begin{eqnarray*}
W_{{\cal C}}^{{\rm R}} (x, y)&=& \sum_{i=0}^n A_i({\cal C}) x^{n-i}y^i\\
&=&\sum_{R \in \Sigma(E)}A_{{\cal C}}(R)x^{n-\dim R}y^{\dim R}\\
&=&\sum_{R \in \Sigma(E)}\left( \sum_{T \in \Sigma(R)}(-1)^{\dim R-\dim T} q^{{\dim R - \dim T \choose 2}}B_{{\cal C}}(T)\right)x^{n-\dim R}y^{\dim R}\\
&=&\sum_{T \in \Sigma(E)}B_{{\cal C}}(T)\left(\sum_{\substack{T \leq R \\ R \in \Sigma(E)}}(-1)^{\dim R-\dim T} q^{{\dim R - \dim T \choose 2}}\right)x^{n-\dim R}y^{\dim R}\\
&=&\sum_{T \in \Sigma(E)}B_{{\cal C}}(T)\sum_{J \leq T^{\perp}}(-1)^{\dim J} q^{{\dim J \choose 2}}x^{n-\dim J-\dim T}y^{\dim J+\dim T}\\
&=&\sum_{T \in \Sigma(E)}B_{{\cal C}}(T)\sum_{J \in \Sigma(T^{\perp})}(-1)^{\dim J} q^{{\dim J \choose 2}}x^{\dim T^{\perp}-\dim J}y^{\dim J+\dim T}\\
&=&\sum_{T \in \Sigma(E)}B_{{\cal C}}(T)y^{\dim T}\prod_{j=0}^{\dim T^{\perp}-1}(x-q^jy)\\
&=&\sum_{T' \in \Sigma(E)}B_{{\cal C}}((T')^{\perp})y^{n-\dim T'}\prod_{j=0}^{\dim T'-1}(x-q^jy)\\
&=&\sum_{T' \in \Sigma(E)}|{\cal C}((T')^{\perp})|y^{n-\dim T'}\prod_{j=0}^{\dim T'-1}(x-q^jy)\\
&=&\sum_{T \in \Sigma(E)}q^{\dim_{{\mathbb F}_q} {\cal C}(T^{\perp})}y^{n-\dim T}\prod_{j=0}^{\dim T-1}(x-q^jy)\\
&=&\sum_{T \in \Sigma(E)}q^{\dim_{{\mathbb F}_q} {\cal C}-\rho (T)}y^{n-\dim T}\prod_{j=0}^{\dim T-1}(x-q^jy)\\
&=&y^{n-\rho (E)/m}\sum_{T \in \Sigma(E)}\left(\frac{1}{y^{1/m}}\right)^{m \dim T-\rho (T)} \left( qy^{1/m}\right)^{\rho (E)-\rho(T)}\prod_{j=0}^{\dim T-1}(x-q^jy)\\
&=&y^{n-\dim_{{\mathbb F}_q} {\cal C}/m}\sum_{T \in \Sigma(E)}f_{P_{{\cal C}}}^T\left(qy^{1/m}, \frac{1}{y^{1/m}}\right)g^{\dim T}(x, y).
\end{eqnarray*}
\qed

\section{Application to the MacWilliams Identity}

In this section, we shall present a MacWilliams type identity for Delsarte rank-metric codes as an application of Theorem \ref{thm:greene-type-id}.

\begin{prop}
\label{prop:dual-enumerator}
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
Then 
$$
W_{{\cal C}^{\perp}}^{{\rm R}} (x, y)=\frac{1}{|{\cal C}|}
\sum_{S \in \Sigma(E)}A_{{\cal C}}(S)\sum_{j=0}^n \sum_{l=0}^{j} {n-\dim S \brack j-l}_q {n-j+l \brack l}_q  (-1)^l q^{{l \choose 2}} q^{m(j-l)}  y^j x^{n-j},
$$
where ${a \brack b}_q$ denote the Gaussian binomial coefficient (or $q$-binomial coefficient) of non-nagative integers $a$ and $b$.
\end{prop}
\pf
By combining Theorem \ref{thm:greene-type-id}, Proposition \ref{prop:duality} and Theorem \ref{thm:duality}, we have that
\begin{eqnarray*}
W_{{\cal C}^{\perp}}^{{\rm R}} (x, y)&=&y^{n-\dim_{{\mathbb F}_q} {\cal C}^{\perp} / m}
\sum_{D \in \Sigma(E)}f_{P_{{\cal C}^{\perp}}}^D\left(qy^{1/m}, \frac{1}{y^{1/m}}\right)g^{\dim D}(x, y)\\
&=&y^{\dim_{{\mathbb F}_q} {\cal C} / m}
\sum_{D \in \Sigma(E)}f_{P_{{\cal C}}}^D\left(\frac{1}{y^{1/m}}, qy^{1/m}\right)g^{n-\dim D}(x, y)\\
&=&y^{\dim_{{\mathbb F}_q} {\cal C} / m}
\sum_{D \in \Sigma(E)} \left( qy^{1/m}\right)^{m \dim D-\rho (D)} \left(\frac{1}{y^{1/m}}\right)^{\rho (E)-\rho(D)}\prod_{j=0}^{n- \dim D-1}(x-q^jy)\\
&=&q^{-\dim_{{\mathbb F}_q} {\cal C}}
\sum_{D \in \Sigma(E)} q^{\dim_{{\mathbb F}_q} {\cal C}-\rho(D)} \left( q^my\right)^{\dim D} \prod_{j=0}^{n- \dim D-1}(x-q^jy)\\
&=& \frac{1}{|{\cal C}|}
\sum_{D' \in \Sigma(E)} q^{\dim {\cal C}(D')} \left( q^my\right)^{n-\dim D'} \prod_{j=0}^{\dim D'-1}(x-q^jy)\\
&=&\frac{1}{|{\cal C}|}
\sum_{D \in \Sigma(E)}B_{{\cal C}}(D) (q^my)^{n-\dim D} \prod_{j=0}^{\dim D-1}(x-q^jy)
\\
&=&\frac{1}{|{\cal C}|}
\sum_{D \in \Sigma(E)}\left(\sum_{S \in \Sigma(D)} A_{{\cal C}}(S)\right) (q^my)^{n-\dim D} \sum_{u=0}^{\dim D} {\dim D \brack u}_q (-1)^u q^{{u \choose 2}}y^u x^{\dim D-u}\\
&=& \frac{1}{|{\cal C}|}
\sum_{S \in \Sigma(E)}A_{{\cal C}}(S)\sum_{j=0}^n {n-\dim S \brack n-j}_q q^{m(n-j)} \sum_{l=0}^{j} {j \brack l}_q (-1)^l q^{{l \choose 2}}y^{n-j+l} x^{j-l}\\
&=& \frac{1}{|{\cal C}|}
\sum_{S \in \Sigma(E)}A_{{\cal C}}(S)\sum_{j=0}^n \sum_{l=0}^{j} {n-\dim S \brack j-l}_q {n-j+l \brack l}_q  (-1)^l q^{{l \choose 2}} q^{m(j-l)}  y^j x^{n-j}.
\end{eqnarray*}
\qed

The equation in Proposition \ref{prop:dual-enumerator} can be viewed as a kind of MacWilliams type identities for Delsarte rank-matric codes.
Now we prove the equivalence between the above equation and the MacWilliams type identity for Gabidulin rank-metric codes proposed in \cite{gadouleau2007, gadouleau2008} by using the concepts of $q$-products and $q$-derivative for homogeneous polynomials.
 
\begin{defi}
\label{defi:q-product}
{\rm
(\cite{gadouleau2007, gadouleau2008})
Let $a(x,y;m)=\sum_{i=0}^r a_i(m)x^{r-i}y^i$ and $b(x,y;m)=\sum_{j=0}^s b_j(m)x^{s-j}y^j$ be two homogeneous polynomials in $x$ and $y$ of degree $r$ and $s$ respectively with coefficients $a_i(m)$ and $b_j(m)$ for $i, j \geq 0$ in turn are real functions of $m$, and are assumed to be zero unless otherwise specified.
The $q$-{\em product} $c(x,y;m)$ of $a(x,y;m)$ and $b(x,y;m)$ is defined to be the homogeneous polynomial of degree $(r+s)$ as follows:
$$
c(x,y;m):=a(x,y;m)*b(x,y;m)=\sum_{u=0}^{r+s}c_u(m)x^{r+s-u}y^u,
$$
where
$$
c_u(m):=\sum_{i=0}^uq^{is}a_i(m)b_{u-i}(m-i).
$$
For $n\geq 0$, the {\em $n$-th $q$-power} of $a(x,y;m)$ is defined recursively:
\begin{eqnarray*}
a(x,y;m)^{[0]}&:=&1,\: \mbox{\rm and}\\
a(x,y;m)^{[n]}&:=&a(x,y;m)^{[n-1]}*a(x,y;m),\: n=1,2,\ldots,
\end{eqnarray*}
}
\end{defi}

\begin{lem}
\label{lem:q-products}
For any $n \geq 0$ and $(0 \leq) l \leq n$,
\begin{enumerate}
\item[{\rm (1)}] $(q^m y)^{[n]}=(q^my)^n$,
\item[{\rm (2)}] $(x-y)^{[l]}*(q^my)^{[n-l]}=(x-y)^{[l]}\times (q^my)^{n-l}$.
\end{enumerate}
\end{lem}
\pf
(1) We shall prove this equation by induction on $n$.
When $n=0$, it follows that $(q^m y)^{[0]}=1=(q^my)^0$.
Suppose that the equation is true for $n-1$, that is,
$(q^m y)^{[n-1]}=(q^my)^{n-1}$.
Then, from the definition of $q$-products, we have that
\begin{eqnarray*}
(q^m y)^{[n]}&=&(q^m y)^{[n-1]}*(q^m y)=(q^my)^{n-1}*q^m y\\
&=&(q^{n-1}\times q^{m(n-1)}\times q^{m-n+1})y^n=(q^my)^n.
\end{eqnarray*}
(2) From Lemma 1 in \cite{gadouleau2007}, it follows that
$$
(x-y)^{[l]}=\sum_{u=0}^l {l \brack u}_q (-1)^uq^{{u \choose 2}}x^{l-u}y^u,
$$
where
$$
{l \brack u}_q:=\prod_{i=0}^{u-1}\frac{1-q^{l-i}}{1-q^{i+1}}.
$$
Then we have that
\begin{eqnarray*}
(x-y)^{[l]}*(q^my)^{[n-l]}&=&\left(\sum_{u=0}^l {l \brack u}_q (-1)^uq^{{u \choose 2}}x^{l-u}y^u\right)*\left(q^{m(n-l)}y^{n-l}\right)\\
&=&\sum_{u=0}^n c_u(m) x^{n-u}y^u,
\end{eqnarray*}
where
\begin{eqnarray*}
c_u(m)&=&0,\: u=0,1,\ldots,n-l-1,\\
c_{n-l+j}(m)&=&q^{m(n-l)}{l \brack j}_q (-1)^jq^{{j \choose 2}},\: j=0,1,\ldots,l.
\end{eqnarray*}
Therefore it follows that
\begin{eqnarray*}
(x-y)^{[l]}*(q^my)^{[n-l]}&=&q^{m(n-l)}y^{n-l}\times \sum_{u=0}^l {l \brack u}_q (-1)^uq^{{u \choose 2}}x^{l-u}y^u\\
&=&(q^my)^{n-l}\times (x-y)^{[l]}.
\end{eqnarray*}
\qed

\begin{defi}
\label{defi:q-transform}
{\rm
(\cite{gadouleau2007, gadouleau2008})
The $q$-{\em transform} of a homogeneous polynomial\\ $a(x,y;m)=\sum_{i=0}^r a_i(m)x^{r-i}y^i$ is defined by
the homogeneous polynomial
$$
\overline{a}(x, y; m):=\sum_{i=0}^r a_i(m) y^{[i]}*x^{[r-i]}.
$$
}
\end{defi}

The following lemma is essential:

\begin{lem}
\label{lem:equations}
\begin{enumerate}
\item[{\rm (1)}] $\displaystyle
{a \brack b}_q={a-1 \brack b}_q+q^{a-b}{a-1 \brack b-1}_q=q^b{a-1 \brack b}_q+{a-1 \brack b-1}_q;
$
\item[{\rm (2)}] $\displaystyle
{a \brack b}_q {b \brack c}_q={a \brack b-c}_q {a-b+c \brack c}_q;
$
\item[{\rm (3)}] $\displaystyle
{a \brack b}_q=\sum_{i=0}^nq^{i(a-b-n+i)}{n \brack i}_q {a-n \brack b-i}_q,\:\:\:n=0,1,\ldots, a;
$
\item[{\rm (4)}] $\displaystyle
{a+b \choose 2}={a \choose 2}+ab+{b \choose 2}.
$
\end{enumerate}
\end{lem}

By combining Proposition \ref{prop:dual-enumerator} and Lemmas \ref{lem:q-products} and \ref{lem:equations}, we have the following MacWilliams type identity for a rank weight enumerator of a Delsarte rank-metric code.

\begin{thm}
\label{thm:MacWilliams}
Let ${\cal C}$ be a Delsarte rank-metric code in $\mbox{\rm Mat}(n\times m, {\mathbb F}_q)$.
Then 
$$
W_{{\cal C}^{\perp}}^{{\rm R}} (x, y)=\frac{1}{|{\cal C}|}\overline{W_{{\cal C}}^{{\rm R}}} (x+(q^m-1)y, x-y).
$$
\end{thm}
\pf
By Corollary 1 in \cite{gadouleau2008}, we have that
$$
(x-y)^{[i]}*(x+(q^m-1)y)^{[n-i]}=\sum_{j=0}^nP_j(i ; m, n)y^jx^{n-j},
$$
where
$$
P_j(i ; m, n):=\sum_{l=0}^j {l \brack j}_q {n-i \brack j-l}_q (-1)^l q^{{l \choose 2}} q^{l(n-i)} \prod_{u=0}^{j-l-1} (q^{m-l}-q^u).
$$
In addition, from Proposition \ref{prop:dual-enumerator}, it is sufficient to prove that
$$
 \sum_{l=0}^{j} {n-i \brack j-l}_q {n-j+l \brack l}_q  (-1)^l q^{{l \choose 2}} q^{m(j-l)}=P_j(i ; m, n).
$$

By using Lemma \ref{lem:equations}, we have that
\begin{eqnarray*}
\mbox{\rm L.H.S.}&=&
 \sum_{l=0}^{j} {n-i \brack j-l}_q \left\{ \sum_{u=0}^i q^{u(n-j-i+u)} {i \brack u}_q {n-j+l-i \brack l-u}_q\right\}  (-1)^l q^{{l \choose 2}} q^{m(j-l)}\\
 &=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-j-i+u)}\sum_{l=0}^{j} {n-i \brack j-l}_q {n-j+l-i \brack l-u}_q  (-1)^l q^{{l \choose 2}} q^{m(j-l)}\\
&=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-j-i+u)}{n-i \brack j-u}_q \sum_{l=0}^{j}  {j-u \brack l-u}_q  (-1)^l q^{{l \choose 2}} q^{m(j-l)}\\
&=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-j-i+u)}{n-i \brack j-u}_q \sum_{l=0}^{j-u}  {j-u \brack l}_q  (-1)^{l+u} q^{{l+u \choose 2}} q^{m(j-l-u)}\\
&=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-j-i+u)}(-1)^u q^{{u \choose 2}}{n-i \brack j-u}_q \sum_{l=0}^{j-u}  {j-u \brack l}_q  (-1)^{l} q^{{l \choose 2}} (q^{m})^{(j-u)^l} (q^u)^l\\
&=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-j-i+u)}(-1)^u q^{{u \choose 2}}{n-i \brack j-u}_q \prod_{i=0}^{j-u-1} (q^m-q^{i+u})\\
&=&
\sum_{u=0}^i {i \brack u}_q q^{u(n-i)}(-1)^u q^{{u \choose 2}}{n-i \brack j-u}_q \prod_{i=0}^{j-u-1} (q^{m-u}-q^{i})\\
&=& P_j(i ; m, n).
\end{eqnarray*}
\qed

\section*{Acknowledgement}
The author would like to thank Prof. Frederique Oggier for introducing him to the concept of rank-metric codes.

\end{document}